# Adaptive Informed Deep Neural Networks for Power Flow Analysis

### v1

@@ V-C Model Sensitivity Analysis
The architecture of PINN4PF and MLP, including the number of hidden layers and neurons per hidden layer, as well as hyperparameters, including the learning rate, the weight decay rate, and the percentage of dropout, are determined by sensitivity analysis to ensure a fair comparison [^1] An exhaustive search across 2000 unique hyperparameter combinations is conducted using the Optuna Python package [40].. The maximum weight for the share of the supervised penalty term ($\beta_{0}$) in (17) is also fine-tuned for PINN4PF. It should be noted that for the MLP, $\beta_{0}=1$ and $\beta_{1}=0$. Sensitivity analysis is not performed for the LR since it involves linear combinations and lacks hyperparameters. The optimal configurations are identified based on minimizing both training and testing MSE, but also ensuring that they remain close enough to indicate generalization ability. All experiments use the 15-bus test system and a noisy dataset comprising $512$ data points.
@@ V-C3 Impact of modifications on the performance of PINN4P
We evaluate the following configurations: (i) a double-head feed-forward NN with ReLU (9) and supervised penalty term (ReLU&Supervised), (ii) a double-head feed-forward NN with adaptive ReLU (10) and supervised penalty term (AdaptiveReLU&Supervised), (iii) a double-head feed-forward NN with ReLU (9) and physical penalty term (ReLU&Physical), and (iv) a double-head feed-forward NN with adaptive ReLU (10) and physical penalty term (AdaptiveReLU&Physical) that is the selected configuration for PINN4PF. Table IV presents a comparative analysis of their performance in terms of direct physical quantities.
@@ VI Discussion
  • Although only the diagonal elements of the admittance matrix are used to develop the loss function, including a physical penalty term imposes additional calculations during training, which makes PINN4PF computationally more expensive and highlights a trade-off between accuracy and computational cost.
@@
  • In modern power systems, it is expected to have a lot of missing data or no measurements. Traditional PF solvers, such as NR, could not converge to high-quality solutions. Nevertheless, this study demonstrates that PINN4PF is not only more robust to noisy data compared to class. Still, it can also learn the complex dynamics of power systems from fewer training samples. These findings highlight the importance of physics-informed components in power system research.
@@ VII Conclusion
This study presents an end-to-end architecture for deep learning-based power flow (PF) analysis called PINN4PF. PINN4PF includes three significant modifications: a double-head feed-forward neural network, an adaptive activation function, and a physics-based loss function. PINN4PF, therefore, offers a straightforward yet effective approach to capturing the complexities inherent in large-scale modern power systems. The application of PINN4PF to four test power systems, including 4-bus, 15-bus, 290-bus, and 2224-bus systems, rigorously evaluates its performance against two baselines: a linear regression model (LR) and a black-box neural network (MLP). The results highlight PINN4PF’s exceptional generalization ability, robustness against noise, data efficiency, and scalability to large-scale power systems. Specifically, PINN4PF consistently achieves lower mean squared errors for direct and derived physical quantities, proving its effectiveness and reliability. In addition, as the test system size increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. This advancement is crucial for enhancing power system operations, especially under extreme conditions and unexpected deviations, making PINN4PF a promising solution for future power flow analysis and optimization tasks.

### v2

@@ V-C Model Sensitivity Analysis
The architecture of PINN4PF and MLP, including the number of hidden layers and neurons per hidden layer, as well as hyperparameters, including the learning rate, the weight decay rate, and the percentage of dropout, are determined by sensitivity analysis to ensure a fair comparison [^1] An exhaustive search across 2000 unique hyperparameter combinations is conducted using the Optuna Python package [48].. The maximum weight for the share of the supervised penalty term ($\beta_{0}$) in (12) is also fine-tuned for PINN4PF. It should be noted that for MLP, $\beta_{0}=1$ and $\beta_{1}=0$. Sensitivity analysis is not performed for LR since it involves linear combinations and lacks hyperparameters. The optimal configurations are identified based on minimizing both training and testing MSE, but also ensuring that they remain close enough to indicate generalization ability. All experiments use the 15-bus test system and a noisy dataset comprising $512$ data points.
@@ V-C3 Impact of modifications on the performance of PINN4P
We evaluate the following configurations: (i) a double-head feed-forward NN with ReLU (5a) and supervised penalty term (ReLU&Supervised), (ii) a double-head feed-forward NN with adaptive ReLU (5b) and supervised penalty term (AdaptiveReLU&Supervised), (iii) a double-head feed-forward NN with ReLU (5a) and physical penalty term (ReLU&Physical), and (iv) a double-head feed-forward NN with adaptive ReLU (5b) and physical penalty term (AdaptiveReLU&Physical) that is the selected configuration for PINN4PF. Table IV presents a comparative analysis of their performance in terms of direct physical quantities. The results show that combining the physics-informed loss function with the adaptive activation function yields superior performance of PINN4PF than the other combinations.
@@ VI Discussion
  • Although only the diagonal elements of the admittance matrix are used to develop the loss function, including a physical penalty term imposes additional calculations during training, which makes PINN4PF computationally more expensive than supervised models and highlights a trade-off between accuracy and computational cost.
@@
  • PINN4PF is trained on different operating conditions, and hence, can generalize to provide approximate solutions even in ill-conditioned scenarios, where traditional methods fail to converge. Future work should explore the use of these approximate solutions as initial guesses (warm starts) for traditional methods and assess their potential to enhance convergence in such challenging cases.
  • In modern power systems, missing or noisy measurements and incomplete topology data are common, often causing traditional PF solvers to fail. PINN4PF addresses these challenges by integrating partial topology information, specifically, the diagonal elements of the admittance matrix, alongside combining supervised and physics-based penalty terms. This design enables PINN4PF to learn complex system dynamics from limited data and enhances robustness against noise or incompleteness in both training samples and topology information. Future work may further assess PINN4PF’s resilience against noisy data in the admittance matrix.
@@ VII Conclusion
This study presents an end-to-end architecture for deep learning-based power flow (PF) analysis called PINN4PF. PINN4PF includes three significant modifications: a double-head feed-forward neural network, an adaptive activation function, and a physics-based loss function. PINN4PF, therefore, offers a straightforward yet effective approach to capturing the complexities inherent in large-scale modern power systems. The application of PINN4PF to four test power systems, including 4-bus, 15-bus, 290-bus, and 2224-bus systems, rigorously evaluates its performance against two baselines: a linear regression model (LR) and a black-box neural network (MLP).
The results highlight PINN4PF’s exceptional generalization ability, robustness against noise, data efficiency, and scalability to large-scale power systems. Specifically, PINN4PF consistently achieves lower mean squared errors for direct and derived physical quantities, proving its effectiveness and reliability. In addition, as the test system size increases, the performance gap between PINN4PF, MLP, and LR becomes more pronounced. This advancement is crucial for enhancing power system operations, especially under extreme conditions and unexpected deviations, making PINN4PF a promising solution for future PF analysis and optimization tasks.